OpenStudy (anonymous):

Factor completely. 9x^2-6xy+y^2-81

OpenStudy (anonymous):

9x² - 6xy + y² - 81 = ( 9x² - 6xy + y² ) - 81 = ( 9x² - 6xy + y² ) - ( 9 )² First bracket = a² - 2ab + b² = ( a - b )² = [ ( 3x )² - 2( 3x )( y ) + ( y )² ] - ( 9 )² = ( 3x - y )² - ( 9 )² Difference of 2 squares = a² - b² = ( a + b )( a - b ) = ( 3x - 9y + 9 )( 3x - 9y - 9 ) Hope this helps

OpenStudy (anonymous):

whts the answer

OpenStudy (anonymous):

(3x - y - 9)(3x - y + 9)